1) Solve the following equations.
a) √X - 1= 3
Answer:
Show work in this space.



b) √x^3 =8
Answer:
Show work in this space.



c) 3^√x^2 =4
Answer:
Show work in this space.

Answer by pradhyumna.kumar@gmail.com(62)

1) Solve the following equations.
a) √X - 1= 3
Answer: √X = 4 => x=16


b) √x^3 =8
Answer: √x^3 =8 => squaring on both sides then u get x^3 = 64 => x = 4



c) 3^√x^2 =4
Answer: 3^√x^2 =4 => i did not understood the problem
may u repete and check the question